The following selected account balances were taken from XYZ Company's general
ledgers for 2022:

                      January 1, 2022       December 31, 2022
  Inventory                  47,000                56,000
  Accounts payable           50,000                28,000
  Salaries payable            4,000                 9,000
  Investments                69,000                75,000
  Accounts receivable        68,000                35,000
  Land                       58,000                88,000
  Notes payable             120,000                95,000
  Unearned revenue           17,000                25,000
  Common stock              100,000               170,000
  Retained earnings          23,000                35,000

The following information was taken from XYZ Company's 2022 income statement:

  Sales revenue                   $420,000
  Cost of goods sold               300,000
  Salaries expense                  88,000
  Net income                      $ 32,000

A) Calculate the amount of cash paid to employees for salaries during 2022.
B) Calculate the amount of cash paid to suppliers for purchases of inventory for 2022.
C) Calculate the amount of cash collected from customers during 2022.
Answer #1

A)Cash paid to employees for salaries = Beginning salaries payable +salaries expense -ending salaries payable

                     = 4000 +88000 - 9000

                      = 83000

b)Cash paid to suppliers for purchase of inventory=Beginning accounts payable +purchase -ending accounts payable

          = 50000+309000 - 28000

             = 331000

Working:

cost of goods sold =Beginning inventory +purchase -ending inventory

     300000 = 47000 + Purchase - 56000

    purchase = 300000+56000-47000

               = 309000

c)cash collected from customers = Beginning accounts receivable+ sales -ending accounts receivable

            = 68000 +420000-35000

          = 453000